A 36 g block of ice is cooled to −77 ◦C. It
is added to 589 g of water in an 74 g copper
calorimeter at a temperature of 26◦C.
Find the final temperature. The specific
heat of copper is 387 J/kg ·
◦C and of ice is
2090 J/kg · ◦C . The latent heat of fusion of
water is 3.33 × 105
J/kg and its specific heat
is 4186 J/kg · ◦C .
Answer in units of ◦C.
1
Expert's answer
2020-04-08T10:27:13-0400

m1= 36 g=0.036 kg - mass of ice

T1= −77 ◦C. - temperature of ice

m2=589 g=0.589 kg - mass of water

C1=2090 J/kg · ◦C - specific heat of ice

"\lambda" = 3.33 × 10^5 J/kg - latent heat of fusion of water

T2=26◦C. - temperature of water

C2= 4186 J/kg · ◦C . - - specific heat of water

m3=74 g=0.074kg - mass of copper

T3=26◦C. - temperature of copper

C3=387 J/kg ·◦C - specific heat of copper is and of ice is

T - ? - final temperature



"m_1C_1(0-T_1) - ice\ heating\ to \ 0^oC\\
m1\lambda - ice\ melting\\
m_1C_2(T-0) - melted\ water\ heating\ \\
m_2C_2(T-T_2) - water\ cooling\\
m_3C_3(T-T_3) - copper\ cooling"


"m_1C_1(0-T_1)+m1\lambda+m_1C_2(T-0)+m_2C_2(T-T_2)+m_3C_3(T-T_3)=0\\"

"0.036 \u00b7 2090 \u00b7 77+0.036 \u00b7 3.33 \u00b7 10^5+0.036 \u00b7 4186 \u00b7 T+0.589 \u00b7 2090\u00b7 (T-26)+0.074\u00b7387\u00b7 (T-26) =0"


"1410.344 · T=14969.368\\
T=10.61^oC"

Answer: "T=10.61^oC"